OpenStudy (anonymous):

Solve (1/x) + x =(1/2) Any help? I'm not sure how to solve this.

OpenStudy (tkhunny):

Write down that x cannot be zero, then multiply by x.

OpenStudy (anonymous):

Lol. Hero is a 99. He gave you the correct answer.

OpenStudy (anonymous):

Can you give me step by step directions, please?

OpenStudy (anonymous):

Hero gave you step by step directions.

OpenStudy (tkhunny):

Or, you can just think about it, instead, x and 1/x are reciprocals. Example. 4 and 1/4 or 6 and 1/6 Funny thing about reciprocals, ONE of them MUST be AT LEAST 1. So, you add up something that is at least 1, plus a little more, it is unlikely that you will find 1/2 in the mix. x + 1/x >= 1

OpenStudy (tkhunny):

Unless, of course, you can use complex numbers. Then we're in business.

OpenStudy (anonymous):

So like (1/x) and x = (x/1)?

OpenStudy (anonymous):

@tkhunny He isn't in that high of a math level. Lol.

OpenStudy (anonymous):

You'd confuse him so much with complex numbers/solutions.

OpenStudy (anonymous):

Are you calling me stupid?

OpenStudy (anonymous):

No, not at all. The level of math you are doesn't include complex numbers/solutions.

OpenStudy (tkhunny):

Try it out. Pick some numbers 1 2 3 4 5 6 7 Find their reciprocals 1 1/1 2 1/2 3 1/3 4 1/4 5 1/5 6 1/6 7 1/7 Add the two 1 + 1/1 = 2 2 + 1/2 = 5/2 = 2.5 3 + 1/3 = 10/3 = 3.3333... 4 + 1/4 = 17/4 = 4.25 5 + 1/5 = 26/5 = 5.2 6 + 1/6 = 37/6 = 6.166666... 7 + 1/7 = 50/7 = 7.1428571... We didn't get anywhere near 1/2 or eve 1, did we?

OpenStudy (anonymous):

1=(1/1)+ (1/1) =(2/2)?

OpenStudy (tkhunny):

That makes no sense. 1 is not equal to 2, no matter what you do in the middle. Just a number and its reciprocal. That's all you get.

OpenStudy (anonymous):

1 =1/1 sorry. So if you add 1 + (1/1) = (2/2)?

OpenStudy (tkhunny):

No, if youadd 1 + 1/1, you get 2/1 = 2,not 1.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

So 2 would be (1/2) and not (2/1)?

OpenStudy (tkhunny):

A positive number and it's reciprocal. One must be greater than 1 and the other must be less than 1 (except for 1, when they are the same.) 2 has reciprocal 1/2 Adding 2 and 1/2 is like this \(2 + \dfrac{1}{2} = \dfrac{4}{2} + \dfrac{1}{2} = \dfrac{4+1}{2} = \dfrac{5}{2} = 2.5\) Common Denominator stuff, right?

OpenStudy (tkhunny):

2 IS 2/1 2's reciprocal is 1/2

OpenStudy (anonymous):

Okay. Sorry. I wasn't thinking right.

OpenStudy (anonymous):

I understand now.

OpenStudy (tkhunny):

Solution to the problem? x + 1/x = 1/2 Algebra and prove that it has No Real Solution. Hero did a good job showing you this, earlier. Logic and demonstrate that there can be no such solution in the Real Numbers. Unique solutions don't care how you find them.

OpenStudy (anonymous):

(1/x) + (x/x) = (1x/x^2)?

OpenStudy (anonymous):

Or is that wrong?

OpenStudy (tkhunny):

\(\dfrac{1}{x} + x = \dfrac{1}{2}\) If we really want to find a common denominator. \(\dfrac{1}{x} + \dfrac{x^{2}}{x} = \dfrac{x}{2x}\) Almost. One more try. \(\dfrac{2}{2x} + \dfrac{2x^{2}}{2x} = \dfrac{x}{2x}\) There it is. Take a close look at your middle term. You started with x and changed it to x/x. Since x/x = 1, this is a very different expression. That is no good. If you were MULTIPLYING the fractions, \(\dfrac{1}{x}\cdot\dfrac{x}{x} = \dfrac{x}{x^{2}}\), but that's a whole different planet.

OpenStudy (anonymous):

So it is better to multiply it?

OpenStudy (tkhunny):

Better? Who cares! What does the problem require? It is not a good question. \(\dfrac{1}{x} + \dfrac{x}{x} = \dfrac{1+x}{x}\) - The denominator doesn't change on addition.

OpenStudy (anonymous):